OpenStudy (anonymous):

Medal For Help :) Which of (3,5), (4,6), (5,7), and (6,8) are solutions to y = x + 2? A. (3,5), and (4,6) B. (4,6), and (5,7)*** My Answer C. none D. all

OpenStudy (anonymous):

@kropot72

OpenStudy (imstuck):

Any ideas of where to start with this? When you find out what to do to solve this and how easy it is, you will smile!

OpenStudy (anonymous):

Can you start me off :)

OpenStudy (imstuck):

In your coordinates above, the first number inside the parenthesis is the x coordinate, and the second is the y. I know you know that; I'm just tying it all in so in the end, you have a clear and complete understanding.

OpenStudy (imstuck):

The idea here is to have whatever your y coordinate is = the x coordinate plus 2. So, for example, if you had a coordinate point of ( 2, 4), you know that y is 4 and x is 2, and filling them into your equation of y = x + 2, you would get 4 = 2 + 2. And 4 does in fact equal 2 + 2 which is 4. See that?

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

so we need to test your points, each one of them, in the equation y = 2 + x to see which ones make the equation "true". Let's look at the first round of choices...

OpenStudy (imstuck):

(3, 5) and (4, 6). Filling in our x's and y's into our equation y = x + 2, using the first set of coordinates we get 5 = 3 + 2. And indeed 5 does equal 3 + 2. S0 (3, 5) works.

OpenStudy (imstuck):

Now let's test the (4, 6).

OpenStudy (imstuck):

6 = 4 + 2. Does 4 + 2 = 6? Yes, it does! So both of those work in that equation! Let's try the second set of choices, just to be sure there isn't more than one choice for your answers, ok?

OpenStudy (anonymous):

D all

OpenStudy (imstuck):

(4, 6) and (5, 7), which you said are your answers. Let's see! We know that (4, 6) works, we just did that one from the first choice. Let's try (5, 7). 7 = 5 + 2. That works too! Wow! We have more than one choice! I have a feeling it is D!

OpenStudy (imstuck):

You're right, because C is "none" and we know it's not none! Good job!

OpenStudy (imstuck):

Thank you for the medal! You just raised my SmartScore 1 point!

OpenStudy (anonymous):

Ty I did all the work on paper with everything you showed me :)